OpenStudy (k8lyn911):

Express the given quantity as a single logarithm. 1/5(ln(x+2)^5)+1/2[ln(x) - ln(x^2+3x+2)^2]

OpenStudy (k8lyn911):

\[\frac{ 1 }{ 5 } \ln (x+2)^5 +\frac{ 1 }{ 2 } \left[ \ln x - \ln (x^2 +3x+2)^2\right]\]

OpenStudy (anonymous):

\[\frac{1}{5}\ln(x+2)^5=\ln(x+2)\] is a good start

OpenStudy (anonymous):

before i combined i would actually distribute the \(\frac{1}{2}\) on the second set of parentheses that would give \[\frac{1}{2}\ln(x)-\ln(x^2+3x+2)\] as the two would cancel

OpenStudy (anonymous):

ok so far?

OpenStudy (k8lyn911):

Yep.

OpenStudy (anonymous):

then \[\frac{1}{2}\ln(x)=\ln(\sqrt{x})\] and then you are good to go

OpenStudy (anonymous):

\[\ln(x+2)+\ln(\sqrt{x})-\ln(x^2+3x+2)\] combine as \[\log(A)+\log(B)-\log(C)=\log(\frac{AB}{C})\]

OpenStudy (anonymous):

oh and one more thing since \(x^2+3x+2=(x+2)(x+1)\) you will be able to cancel a factor of \(x+2\) top and bottom

OpenStudy (k8lyn911):

So the answer is \[\ln \frac{ \sqrt{x} }{ x+1 }\] ?

OpenStudy (anonymous):

that is what i get, yes

OpenStudy (anonymous):

seem reasonable?

OpenStudy (k8lyn911):

Yes. It makes a lot more sense now. Thank you so much! :)

OpenStudy (anonymous):

yw now i have a question

OpenStudy (anonymous):

is that really a chuck e cheese hat?

OpenStudy (k8lyn911):

Yes. Yes, it is.

OpenStudy (anonymous):

ho ho ho nice

OpenStudy (k8lyn911):

I didn't want to get a sunburn on my head, because that makes showering really terrible. I'm not really a hat person, though, so I had to borrow one.

OpenStudy (anonymous):

guess there was a 10 year old handy

OpenStudy (k8lyn911):

There was a five-year old handy, yes. xD